# FLex&Chill: Improving Local Federated Learning Training with Logit Chilling

### v1

@@ Abstract
Federated learning are inherently hampered by data heterogeneity: non-iid distributed training data over local clients. We propose a novel model training approach for federated learning, FLex&\&&Chill, which exploits the Logit Chilling method. Through extensive evaluations, we demonstrate that, in the presence of non-iid data characteristics inherent in federated learning systems, this approach can expedite model convergence and improve inference accuracy. Quantitatively, from our experiments, we observe up to 6$\times$ improvement in the global federated learning model convergence time, and up to 3.37% improvement in inference accuracy.
@@ 1 Introduction
Federated learning represents a paradigm enabling the training of effective models within distributed environments without explicitly exposing local data McMahan et al. (2017). This approach provides opportunities for communication and computation efficiency during training by sharing model parameters/gradients learned independently from local leaning processes while suppressing raw data transmissions. Furthermore, since training operations take place in a de-centralized manner, federated learning opens the possibility to parallelize the model training process via the assistance of local client devices. It presents an attractive solution for handling and exploiting the vast amounts of data generated by many mobile and Internet of Things (IoT) device entities consisting the overall system Park et al. (2023a).
@@
In contrast to conventional centralized training methods, federated learning faces challenges associated with delayed model convergence and limited performance. A fundamental challenge of federated learning arises from the disparity of training data used at each client device. Federated stochastic optimization encounters difficulties in determining optimal parameters, leading to increased communication overhead between local devices and the central server, as well as the need for additional training on (typically resource limited) local devices Zhao et al. (2018).
@@
To address these challenges, a number of prior research have focused on optimizing the local training operations on federated learning clients or on refining the server-side aggregation process. For instance, FedProx Li et al. (2020) controls the number of iterations for each local device, aiming to train models resilient to challenges posed by non-independent and non-iid data environments. SCAFFOLD Karimireddy et al. (2020) achieves expedited convergence and improved model accuracy McMahan et al. (2017) by introducing a correction term during the model aggregation phase to balance the influence of each client. These operations alleviate the problems posed by the non-iid environment, a common dataset configuration for federated learning clients.
@@
This paper proposes a novel model training approach for federated learning, FLex&\&&Chill (Federated Learning EXploiting logit Chilling), that targets to boost the convergence and model performance via temperature scaling, namely logit chilling, during the local model training operations. Note that a typical neural network generates a probability ($q_{i}$) for each class from its neural network logits ($z_{i}$) by applying a softmax operation (Eq. 1) with the temperature $T$ conventionally set to 1. However, it is not uncommon in other application scenarios to exploit even higher temperature values (i.e., $T\in(1,\infty)$), for example, in knowledge distillation Hinton et al. (2015); Touvron et al. (2021) or neural network calibration Guo et al. (2017).
$\xi(z,i,T)=q_{i}=\frac{exp(z_{i}/T)}{\sum_{j}exp(z_{j}/T)}$ (1)
@@
When employing a temperature $T$ higher than 1, the probability distribution across classes naturally undergoes smoothing due to the narrowed interval between the logits Guo et al. (2017), making such an approach suitable for applications such as knowledge distillation. Notably, despite altering the probabilities, the order of labels’ probability remains unaffected with respect to changes in $T$. While these works successfully explore the impact of varying temperatures for neural network operations, they have mainly focused on applying high temperatures after the model training process, mostly for the inference operations. Raising $T$ to a higher value reduces the confidence of the training operations. This is especially useful to tame DNNs which are typically overconfident.
@@
On the other hand, in federated learning, where disparate training data is common, increasing the model confidence of each individually trained client model (using a lower $T$) prior to their parameter aggregation can be a more effective approach. Our investigation involves measuring both the gradient flow and the data position shift within the representation space to provide guidelines in this unexplored territory.
@@
Specifically, this work targets to show that utilizing fractional temperature values (i.e., $T\in(0,1)$) during the training process enables effective gradient propagation towards the input layer when applied to federated learning data distribution scenarios. Consequently, we observed that lower temperatures facilitate more efficient data updates within the representation space when dealing with non-iid datasets, a common characteristic of federated learning systems. From our evaluations with three datasets and three baseline federated learning models, we show that these effects accelerate the federated learning model convergence time by up to $\mathbf{6.00\times}$ and improves the inference accuracy by up to $\mathbf{3.37\%}$, despite potential concerns on model instability when training with with low temperatures. We summarize the contributions of this work in three-fold as the following:
@@ 2.0.1 Temperature scaling
In machine learning, the output probability distribution generated by a model serves diverse purposes. Primarily, classification models offer probabilities for each class, showcasing the model’s confidence in its predictions. This aids in assessing how certain the model is in its predictions (i.e., classification outputs) and allows consideration of the likelihood of each class prior to making a final classification decision. Moreover, this distribution offered by the model helps in understanding the uncertainty of a model on its decisions Pearce et al. (2021); Park et al. (2023b).
@@
In neural networks, computing probabilities from logits often involves employing softmax operations, as shown in Eq. 1. Here, the temperature parameter ($T>0$) controls the probability distribution’s concentration towards one class, and is used to manage how the categorical probability is dispersed over the possible classes Jang et al. (2016). Notably, this categorical probability aligns with the concept of entropy in information theory, offering a perspective where the temperature parameter governs the level of disorder or uncertainty within the distribution. Specifically, when $T\in(1,\infty)$ is applied, the probability distribution tends to be smoother due to increased $T$. This arises from the diminished differences between logit values, enhancing the distribution’s entropy. In contrast, with $T\in(0,1)$ within the softmax function, the probability distribution becomes sharper, since lower temperatures amplify the differences between logits, effectively polarizing the distribution and concentrating probabilities closer to the maximum logit value. Consequently, the model becomes more confident in its predictions, as it allocates higher probabilities to fewer classes, thereby reducing uncertainty in the distribution.
@@
Therefore, temperature scaling operates as a means to adjust the probability distribution within a neural network-based system. Previous work in Knowledge Distillation Hinton et al. (2015); Touvron et al. (2021) exploit the temperature value by applying a high temperature $T\in(1,\infty)$ to achieve a smoothed probability distribution. This manipulation aims to magnify small values within the distribution, often representing hidden information encapsulated within the model. Additionally, in generative models Shih et al. (2023); Wang et al. (2020), temperature scaling regulates the randomness inherent in model outputs. This technique allows for the control of variability levels in generated samples, enabling adjustments to generated output diversity. Temperature scaling is also used in in model calibration Guo et al. (2017). Applying temperature adjustments to the model’s output probabilities, typically through a scaling factor $T$, aids in aligning these probabilities more accurately with the true likelihoods or uncertainties present in the data. Likewise, previous work have focused on applying high temperature values for various purposes, mostly for the purpose of conservatively better-understanding the full dataset without bias.
2.0.2 Federated learning
@@
Despite its benefits, federated learning poses several challenges. A body of work have tackled the issue of communication overhead stemming from the exchange of frequent model parameter updates. A number of previously proposed work have proposed schemes to effectively abstract the model parameters in various forms to relieve the communication and networking overhead Park et al. (2023a). Furthermore, federated learning systems are prone to facing issues rising from data heterogeneity caused from the diversity or differences in the characteristics, distributions, or formats of data across federated learning clients. Previous work have attempted to address this challenge by addressing relevant issues such as variations in feature/input variations Yu et al. (2021) and spatio-temporal variations Liu et al. (2020).
Figure 1: Distribution of gradient norm observed at the input layer for correctly and incorrectly inferenced samples with varying temperature values. We observe noticeable improvements in gradient flow to the when training with lower temperatures.
@@ 3.1 Impact of Training with Lower Temperatures
As aforementioned, conventionally, temperature scaling has primarily been focused on calibrating the probability distribution of a model’s output after the model training phase, with limited exploration into its application during the training process itself. In this section, we present empirical results on three initial experiments specifically crafted to showcase the implications of employing temperature scaling throughout the training phase operations in different dimensions. Through these investigations, our aim is to establish a foundation for comprehending the potential implications and advantages associated with the utilization of temperature scaling techniques in neural network training. We conduct our experiments with the CIFAR10 dataset Krizhevsky et al. (2009) and use a 2-layered CNN architecture McMahan et al. (2017) as the baseline neural network architecture.
@@ 3.1.1 Gradient flow patterns
Our first experiment focuses on assessing the norm of computed gradients within the input layer for various temperature values ($T$), and we present sample distribution plot of how the gradients flow for 50K training samples in Figure 1. Note that here we separate the plots for the correctly inferenced samples (in the foward pass) and the incorrect ones to observe their differences. The propagation of gradients to the input layer serves as an indicative metric for the model’s learning efficacy. Successful gradient propagation to the input layer implies that the model effectively utilizes and comprehends information from diverse features within the input data during the training phase Evci et al. (2022).
As shown in Figure 1, the mode of distribution shifts to larger values with decreasing temperatures for both correct and incorrect predictions. One reason behind this mode shift stems from the gradient function characteristics of softmax. Eq. 2 presents the gradient of cross-entropy for the softmax activation with varying temperature $T$ and hard target, one-hot encoded labels. For brevity, we present the final derivation result and leave details in the Appendix. Since the gradient is multiplied by $\frac{1}{T}>1$, for low temperatures $T<1$, we can expect a gradient boost.
$\frac{\partial\xi}{\partial{z_{i}}}=\frac{1}{T}(p_{i}-y_{i})=\frac{1}{T}(\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}-y_{i})$ (2)
Furthermore, from the perspective of the loss term $(p_{i}-y_{i})$, when the model yields an incorrect decision for a given sample, the loss term significantly increases as $p_{i}\xrightarrow{}0$ (more details later in this section). The double-folded effect accounts for substantial mode shift with thick and long tails in the gradient distribution for lower training temperatures.
@@ 3.1.2 Data position shift in the representation space
For this experiment, we configure an initial neural network and perform inference for 1,000 samples to identify the results’ positions on the representation space. Among the 1,000, we select the ones that were misclassified, and recompute their positions after performing one training operation using that specific data point and compute the difference in positions on the representation space with respect to the decision boundary. By doing so, we can observe how drastically one training operation affects the next inference output with respect to different temperatures used in the training phase.
Figure 2: Distributions depicting differences between distances to the decision boundary before and after model updates for varying training temperatures. Notice that lower temperatures show noticeable shift in estimations’ positions on the representation space, suggesting their aggressiveness in modifying the model even with a small number of training samples.
@@
3.1.3 Impact of temperature on individual samples
Figure 3: Probability distribution for single training sample (class index 4) with different training temperatures. A low $T$ hallucinates the model’s probability distribution, causing the neural network to be trained more aggressively even with a small number of samples.
We now examine the impact of different degrees of temperatures on a single sample prediction. Specifically, we take a look at the probability distribution of prediction for a single sample with varying temperature during training process. In Figure 3 we present the the resulting probability distribution for a single training sample (with ground truth index of 4). As the plots show, a higher temperature, $T$=4 in this case, leads to a widely spread probability distribution, and the neural network tends to less confident in prediction. On the other hand, with lower temperatures the information embedded in the data sample becomes more concentrated to the predicted label. If the prediction is incorrect, this increased prediction on the wrong label results in reduced probability on the correct label and increases the loss. This leads to more aggressive changes despite training from a single sample.
In federated learning, clients will train themselves with a small amount of local data (much smaller than typical centralized training). For the models trained at the clients to hold value at the server during the aggregation process, it is important that the local data characteristics are effectively embedded in its model parameters despite the small local dataset size. The aggressive behavior when applying low training temperatures that we see in Figure 3 and throughout our initial experimental results serve as evidence that the core information from each training element can be effectively extracted and learned at the neural network when low temperatures are applied in the federated learning training process.
@@ 3.2 FLex&\&&Chill and Logit Chilling
Based on such empirical observations, we hypothesise that training neural networks with lower temperatures can benefit the model convergence and performance for systems exploiting with non-iid data, which is common for federated learning scenarios. In FLex&\&&Chill we term this as logit chilling, as a we ‘chill down’ the temperature at the softmax operations, thereby altering the logit computation derived from the neural network during the training process. The focus of a FLex&\&&Chill-based federated learning system is to exploit logit chilling to train a client-side model with low temperature values. As we observed through the studies above and as our evaluations will show, this has an effect of effectively exploiting crucial information from the training samples to actively alter the model towards the new knowledge.

### v2

# Improving Local Training in Federated Learning via Temperature Scaling

@@ Abstract
Federated learning is inherently hampered by data heterogeneity: non-i.i.d. training data over local clients. We propose a novel model training approach for federated learning, FLex&\&&Chill, which exploits the Logit Chilling method. Through extensive evaluations, we demonstrate that, in the presence of non-i.i.d. data characteristics inherent in federated learning systems, this approach can expedite model convergence and improve inference accuracy. Quantitatively, from our experiments, we observe up to 6$\times$ improvement in the global federated learning model convergence time, and up to 3.37% improvement in inference accuracy.
@@ 1 Introduction
Federated learning represents a paradigm enabling the training of effective models within distributed environments without explicitly exposing local data McMahan et al. (2017). This approach provides opportunities for communication and computation efficiency during training by sharing model parameters/gradients learned independently from local learning processes while suppressing raw data transmissions. Furthermore, since training operations take place in a decentralized manner, federated learning opens the possibility of parallelizing the model training process via the assistance of local client devices. It presents an attractive solution for handling and exploiting the vast amounts of data generated by many mobile and Internet of Things (IoT) device entities consisting of the overall system Park et al. (2023a).
@@
In contrast to conventional centralized training methods, federated learning faces challenges associated with delayed model convergence and limited performance. A fundamental challenge of federated learning arises from the disparity of training data used at each client device. Federated stochastic optimization encounters difficulties in determining optimal parameters, leading to increased communication overhead between local devices and the central server, as well as the need for additional training on (typically resource-limited) local devices Zhao et al. (2018).
@@
To address these challenges, a number of prior research have focused on optimizing local training operations on federated learning clients or on refining the server-side aggregation process. For instance, FedProx Li et al. (2020) controls the number of local device iterations, aiming to train models resilient to challenges posed by non-independent and non-i.i.d. data environments. SCAFFOLD Karimireddy et al. (2020) achieves expedited convergence and improved model accuracy McMahan et al. (2017) by introducing a correction term in the model aggregation phase to balance the influence of each client. These operations alleviate challenges posed by the non-i.i.d. environment, a common dataset configuration for federated learning clients.
@@
This paper proposes a novel model training approach for federated learning, FLex&\&&Chill (Federated Learning EXploiting logit Chilling), that targets to boost the convergence and model performance via temperature scaling, namely logit chilling, during the local model training operations. Note that a typical neural network generates a probability ($p_{i}$) for each class from its neural network logits ($z_{i}$) by applying a softmax operation (Eq. 1) with the temperature $T$ conventionally set to 1. However, it is not uncommon in other application scenarios to exploit even higher temperature values (i.e., $T\in(1,\infty)$), for example, in knowledge distillation Hinton et al. (2015); Touvron et al. (2021) or neural network calibration Guo et al. (2017).
$\xi(z,i,T)=p_{i}=\frac{exp(z_{i}/T)}{\sum_{j}exp(z_{j}/T)}$ (1)
@@
When employing a temperature $T~{}>$ 1, the probability distribution across classes undergoes smoothing due to the narrowed interval between logits Guo et al. (2017), making such an approach suitable for applications such as knowledge distillation. Notably, despite altering the probabilities, the order of labels’ probability remains unaffected with respect to changes in $T$. While these works successfully explore the impact of varying temperatures for neural network operations, they mainly focus on applying high temperatures after the model training process, mostly for inference operations. Since increasing $T$ reduces the training confidence, such schemes are effective in taming overconfident DNNs.
@@
On the other hand, in federated learning, where disparate training data is common, increasing model confidence of each individually trained client model (using small $T$) prior to parameter aggregation can be a more effective approach. Our investigation involves measuring both the gradient flow and data position shift within the representation space to provide guidelines in this unexplored territory.
@@
Specifically, this work targets to show that utilizing fractional temperature values (i.e., $T\in(0,1)$) during the training process enables effective gradient propagation towards the input layer when applied to federated learning data distribution scenarios. Consequently, we observed that lower temperatures facilitate more efficient data updates within the representation space when dealing with non-i.i.d. datasets, a common characteristic of federated learning systems. From our evaluations with three datasets and three baseline federated learning models, we show that these effects accelerate the federated learning model convergence time by up to $\mathbf{6.00\times}$ and improve the inference accuracy by up to $\mathbf{3.37\%}$, despite potential concerns on model instability when training with with low temperatures. We summarize the contributions of this work in three-fold as the following:
@@ 2.1 Temperature scaling
In machine learning, a model’s output probability distribution serves diverse purposes. Primarily, classification models offer probabilities for each class, showcasing the model’s confidence in its predictions. This certainty measure allows consideration of the likelihood of each class prior to making a final classification decision. Moreover, this distribution offered by the model helps in understanding the uncertainty of a model on its decisions Park et al. (2023b); Pearce et al. (2021).
@@
For neural networks, computing probabilities from logits often involves employing softmax operations, as in Eq. 1. Here, the temperature parameter ($T>0$) controls the probability distribution’s concentration towards one class, and manages how the categorical probability is dispersed over possible classes Jang et al. (2016). Notably, this categorical probability aligns with the concept of entropy in information theory, offering a perspective where $T$ governs the level of disorder or uncertainty within the distribution. Specifically, with $T\in(1,\infty)$, the probability distribution tends to be smoother as $T$ increases. This arises from the diminished differences between logit values, enhancing the distribution’s entropy. In contrast, $T\in(0,1)$ sharpens the probability distribution, since a lower $T$ amplifies inter-logit differences, effectively polarizing the distribution and concentrating probabilities closer to the maximum logit value. Consequently, model predictions are made more confidently, by allocating higher probabilities to fewer classes, thereby reducing uncertainty in the distribution.
@@
Therefore, temperature scaling operates as a means to adjust the probability distribution within a neural network-based system. Previous work in Knowledge Distillation Hinton et al. (2015); Touvron et al. (2021) exploit the temperature value by applying a high temperature $T\in(1,\infty)$ to achieve a smoothed probability distribution. This manipulation aims to magnify small values within the distribution, often representing hidden information encapsulated within the model. Additionally, in generative models Shih et al. (2023); Wang et al. (2020), temperature scaling regulates the randomness inherent in model outputs. This technique allows for the control of variability levels in generated samples, enabling adjustments to generated output diversity. Temperature scaling is also used in model calibration Guo et al. (2017). Applying temperature adjustments to the model’s output probabilities, typically through a scaling factor $T$, aids in aligning these probabilities more accurately with the true likelihoods or uncertainties present in the data. Likewise, previous work has focused on applying high-temperature values for various purposes, mostly for the purpose of conservatively better understanding the full dataset without bias.
2.2 Federated learning
@@
Despite its benefits, federated learning poses several challenges. A body of work has tackled the issue of communication overhead stemming from the exchange of frequent model parameter updates. A number of previously proposed works have proposed schemes to effectively abstract the model parameters in various forms to relieve the communication and networking overhead Park et al. (2023a). Furthermore, federated learning systems are prone to facing issues arising from data heterogeneity caused by the diversity or differences in the characteristics, distributions, or formats of data across federated learning clients. Previous works have attempted to address this challenge by addressing relevant issues such as variations in feature/input variations Yu et al. (2021) and spatio-temporal variations Liu et al. (2020).
@@ 3.1 Impact of Training with Lower Temperatures
As aforementioned, conventionally, temperature scaling has primarily been focused on calibrating the probability distribution of a model’s output after the model training phase, with a limited exploration into its application during the training process itself. In this section, we present empirical results on three initial experiments specifically crafted to showcase the implications of employing temperature scaling throughout the training phase operations in different dimensions. Through these investigations, our aim is to establish a foundation for comprehending the potential implications and advantages associated with the utilization of temperature scaling techniques in neural network training. We conduct our experiments with the CIFAR10 dataset Krizhevsky et al. (2009) and use a 2-layered CNN architecture McMahan et al. (2017) as the baseline neural network architecture. For experimental details, see APPENDIX B
@@ 3.1.1 Gradient flow patterns
Figure 1: Distribution of gradient norm at input layer for correctly/incorrectly inferred samples with varying training temperatures.
Figure 2: Distributions depicting differences between distances to the decision boundary before and after model updates for varying training temperatures. Notice that lower temperatures show a noticeable shift in estimations’ positions on the representation space, suggesting their aggressiveness in modifying the model even with a small number of training samples.
Our first experiment focuses on assessing the norm of computed gradients within the input layer for various temperature values ($T$), and we present a sample distribution plot of how the gradients flow for 50K training samples in Figure 2. Note that here we separate the plots for the correctly inferred samples (in the forward pass) and the incorrect ones to observe their differences. The propagation of gradients to the input layer serves as an indicative metric for the model’s learning efficacy. Successful gradient propagation to the input layer implies that the model effectively utilizes and comprehends information from diverse features within the input data during the training phase Evci et al. (2022).
As shown in Figure 2, the mode of distribution shifts to larger values with decreasing temperatures for both correct and incorrect predictions. One reason behind this mode shift stems from the gradient function characteristics of softmax. Eq. 2 presents the gradient of cross-entropy for the softmax activation ($\mathcal{L}$) with varying temperature $T$ and hard target, one-hot encoded labels. For brevity, we present the final derivation result and leave details in the Appendix (see Appendix A). Since the gradient is multiplied by $\frac{1}{T}>1$, for low temperatures $T<1$, we can expect a gradient boost.
$\frac{\partial\mathcal{L}}{\partial{z_{i}}}=\frac{1}{T}(p_{i}-y_{i})=\frac{1}{T}(\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}-y_{i})$ (2)
Furthermore, from the perspective of the loss term $(p_{i}-y_{i})$, when the model yields an incorrect decision, the loss term significantly increases as $p_{i}\xrightarrow{}0$. The double-folded effect accounts for a substantial mode shift with thick and long tails in the gradient distribution for lower temperatures.
@@ 3.1.2 Data position shift in the representation space
For this experiment, we configure an initial neural network and perform inference for 1,000 samples to identify the results’ positions on the representation space. Among the 1,000, we select the ones that were misclassified, recompute their positions after performing one training operation using that specific data point, and compute the difference in positions on the representation space concerning the decision boundary. By doing so, we can observe how drastically one training operation affects the next inference output to different temperatures used in the training phase.
@@
3.1.3 Visualization in the representation space
To examine the effect of applying a low temperature during the model training phase, we conducted a simple example that classifies 2-dimensional data via logistic regression. We split the data into three different subsets and assign them as three devices with non-i.i.d data. The Softmax layer was applied at different temperatures ($T$=1.0, 0.5) during the local training phase and average model weights for aggregation after 10 local updates.
Figure 3 illustrates the classification boundaries and samples distributed in the representation space for the three clients and the aggregated model. The results suggest that applying $T$ = 0.5 effectively reflects the local dataset characteristics compared to $T$ = 1.0. Specifically, as an example, for clients 0, 1 and the aggregate model, using $T$ = 1.0 fails to make a clear distinction between the red and light green samples; whereas, when applying $T$ = 0.5, this distinction becomes pronounced.
Figure 3: Example of data points in the 2D representation space with their respective classification boundaries for different federated learning clients with varying training temperatures.
@@ 3.2 FLex&\&&Chill and Logit Chilling
Based on such empirical observations, we hypothesize that training neural networks with lower temperatures can benefit the model convergence and performance for systems exploiting non-i.i.d. data, which is common for federated learning scenarios. In FLex&\&&Chill we term this as logit chilling, as we ‘chill down’ the temperature at the softmax operations, thereby altering the logit computation derived from the neural network during the training process. The focus of a FLex&\&&Chill-based federated learning system is to exploit logit chilling to train a client-side model with low-temperature values. As we observed through the studies above and as our evaluations will show, this has an effect of effectively exploiting crucial information from the training samples to actively alter the model towards the new knowledge.